Standing at a crosswalk, you hear a frequency of 5
Answer #1

Actual frequency f = f

f ' = 520 Hz

We know f ' = [V/(V-v)]f

        520 = [(343)/(343-v)]f       

           f = [(343-v)/343]520         ----------( 1)

f " = 430 Hz

We know f " = [V/(V+v)]f

        430 = [(343)/(343+v)]f       

           f = [(343+v)/343]430         ----------( 2)

From equations(1) and(2),

[(343-v)/343]520   = [(343+v)/343]430

(343-v) 52 = (343+v) 43

1.209 (343-v) = 343+v

414.79 - 1.2093 v = 343 +v

             2.2093 v = 414.79 - 343

                          = 32.5 m/s